A 25.0L tank contains 2.0kg of neon gas with a pressure of 98.0 atm at 28 degrees celcius. How many kilograms of oxygen can fill the tank under the same pressure and temperature?
1
Expert's answer
2017-09-25T09:35:07-0400
pV = nRT
With all other parameters being equal, number of moles must also remain the same. Therefore:
n = m/M
M (Ne) = 20.2 g/mol n (Ne) = 2000/20.2 = 99 mol
M (O 2 ) = 32.0 g/mol m (O 2 ) = n · M = 99 · 32 = 3162 g = 3.162 g